# An Analysis of LIGO Glitches Using t-SNE During the First Part of the Fourth LIGO-Virgo-KAGRA Observing Run

### v1

@@ 1 Introduction
Gravitational waves, predicted by Einstein’s general theory of relativity, have revolutionized our understanding of astrophysical phenomena by providing a new observational window into the Universe. Since the first direct detection of gravitational waves in 2015 by the Laser Interferometer Gravitational-Wave Observatory (LIGO) and the Virgo collaboration [LIGOScientific:2014pky], ground-based detectors have been actively searching for signals during dedicated observing runs. Together with Virgo in Italy [VIRGO:2014yos] and KAGRA in Japan [akutsu2021overview], LIGO forms the LVK collaboration, which brings together a large community of researchers working toward the common goal of detecting an increasing number of gravitational-wave events.
@@
LIGO consists of two observatories designed to detect these spacetime ripples: one located in Hanford, Washington (LHO), and the other in Livingston, Louisiana (LLO). During gravitational wave searches, observatories operate in observing mode, and between observing runs, planned pauses are implemented to upgrade instrumentation and improve sensitivity [capote2025advanced]. The first detection occurred during the first observing run (O1), which lasted from September 12, 2015, to January 2016. On January 16, 2024, LIGO concluded the first part of its fourth observing run (O4a), which had started on May 24, 2023. As registered in the Gravitational-Wave Candidate Event Database (GraceDB) [GraceDB], there are already more than two hundred significant detection candidates identified during O4 produced by the mergers of compact objects such as black holes and neutron stars.
@@
LIGO also records various noise transients arising from instrumental or environmental factors, including ground motion, wind, and other disturbances [davis2021ligo, soni2025ligo]. These non-astrophysical transient signals are commonly referred to as glitches. Glitches are non-stationary artifacts that contribute to the background in gravitational wave searches [glanzer2023data, guimaraes2023bridging, yarbrough2025pinch] and, in some cases, can closely mimic signals from compact binary coalescences (CBC). Efforts within CBC searches to mitigate this challenge are therefore essential and rely on different pipelines, such as PyCBC [usman2016pycbc, dal2021real], GstLAL [messick2017analysis, sachdev2019gstlal], MBTA [adams2016low, aubin2021mbta], and SPIIR [hooper2012summed, chu2017low, chu2022spiir], which incorporate likelihood weighting and other strategies to reduce the impact of glitches on search sensitivity.
@@
Glitch morphologies are typically analyzed using spectrograms [chatterji2004multiresolution], which facilitate their classification into distinct glitch groups. Different tools have been developed to analyze these unwanted signals. BayesWave [cornish2015bayeswave, cornish2021bayeswave] and AWaRe [chatterjee2025no], for instance, employ distinct approaches to reconstruct gravitational-wave signals contaminated by glitches. gengli [lopez2022simulating, lopez2025simulation] is an open-source glitch generator useful to support a variety of testing and validation tasks, while other tools focus on glitch classification, improving classifier performance, or distinguishing true astrophysical signals from spurious noise transients [PhysRevD.97.101501, powell2017classification, chan2024gwskynet, razzano2023gwitchhunters, alvarez2024gspynettree]. One example is Gravity Spy, an interdisciplinary project that integrates machine learning, citizen science, and input from LIGO scientists to classify glitches [zevin2024gravity, zevin2017gravity]. Glitches are named based on their morphology and/or known or suspected physical origin. Examples of their spectrogram morphologies can be found in [zevin2017gravity, glanzer2023data, zoogravityspy, cabero2019blip].
@@
The main motivation behind these efforts is to support the overarching goal of glitch studies: identifying their physical origin to eliminate or mitigate their impact on gravitational-wave detection. This work presents an analysis of the main glitch groups observed during O4a, investigating their potential correlations with environmental and/or instrumental variations using the methodology detailed in [ferreira2024using]. This investigation identifies patterns among glitches using the information provided by Omicron[robinet2016omicron, robinet2020omicron], the primary tool in LIGO used to detect noise transients in detector data. When excess power is identified in the data, Omicron characterizes it by analyzing its time-frequency tiles in spectrograms produced from the $Q$-transform[brown1991calculation]. This information is stored in unclustered files, and after preparing these data, we employ t-distributed Stochastic Neighbor Embedding (t-SNE) [van2008visualizing], a dimensionality reduction technique, to project the high-dimensional output of Omicron (which includes numerous tile-based features) into a two-dimensional space.
@@
As an unsupervised technique, t-SNE does not require prior knowledge or a training dataset. Instead, it relies on probability-based similarity between data points to preserve local relative distances in the lower-dimensional representation. This makes t-SNE particularly useful for uncovering patterns in data without predefined assumptions. The resulting projections often reveal meaningful structures; as shown in [ferreira2024using, ferreira2022comparison], cross-checking the clusters obtained from t-SNE with Gravity Spy labels demonstrates that, when applied to unclustered Omicron information, t-SNE produces groups that closely correspond to known glitch classes.
@@
In this work, the primary objective is to track the temporal evolution of glitches during the O4a observing run. To facilitate this analysis, we first apply t-SNE to project the high-dimensional data into a two-dimensional space and then use Agglomerative Clustering [xu2005survey] to group similar glitches and assign labels. The Silhouette Score [rousseeuw1987silhouettes] is used to evaluate clustering quality and determine the optimal number of groups. These group labels serve primarily as a means to organize and visualize the temporal behavior of distinct glitch populations, to support their possible association with environmental or instrumental conditions, and to enable the classification of previously unseen data. This approach is entirely unsupervised and based solely on features derived from Omicron triggers. Further details about the method can be found in [ferreira2024using].
@@ 2 Glitches during O4a at LLO
The rate of glitches with a signal-to-noise ratio (SNR) exceeding 10 at LLO was higher during O4a (approximately 31 glitches per hour) compared to the third observing run, O3 (approximately 26 glitches per hour)[soni2025ligo]. The glitch rate during O3 was significantly influenced by the prevalence of scattering glitches, which dominated the transient population in that period. Scattering glitches occur when stray light couples back into the main beam after interacting with surfaces, whose identification is particularly challenging. Substantial efforts during O3 successfully identified and mitigated some of these scattering surfaces, resulting in a considerable reduction in the occurrence of such glitches[soni2020reducing, soni2024modeling].
Despite these mitigation efforts, the transient rate during O4a surpassed that of O3; this increase was primarily driven by glitches occurring at low frequencies. Figure 1 illustrates the distribution of glitch frequencies below 100 Hz at LLO during O4a, highlighting this trend. The apparent gap in glitches around $14\text{\,}\mathrm{H}\mathrm{z}$ is due to masking by noise from the test masses’ roll modes and calibration lines at those frequencies. Additional peaks are also visible in the histogram, notably around 32, 39, 49, 61, 75, and 93 Hz. We compare the GravitySpy classes [zevin2017gravity] in each of these frequency bins, and spectrogram analysis shows that the Tomte class accounts for more than 50% of glitches with peak frequencies near 32 and 39 Hz, while the 49 Hz peak consists of approximately 50% Blip Low Frequency. At 61 Hz, the dominant class is Blip Low Frequency; the peak around 75 Hz is primarily associated with Koi Fish, and the 93 Hz region includes a mix of Koi Fish and other classes. Some classes can share similar peak frequencies, but due to the complex nature of glitches, peak frequency alone is usually insufficient to distinguish between certain groups. This underscores the importance of analyzing glitches based on their morphology.
@@
The increase in glitch occurrence at lower frequencies is associated with the improved sensitivity of the interferometer in this frequency range [soni2025ligo]. Figure 2 illustrates this by comparing the characteristic strain at low frequencies (10–30 Hz) during O3, shown in blue, with that of O4a, shown in pink. This comparison highlights the enhanced sensitivity of the interferometer during O4a, which resulted in a higher detection rate of glitches (and of gravitational waves) within this frequency range. The blue dots represent the amplitude of scattering glitches over an eight-hour period on December 27, 2019 (during O3). Since these glitches were associated with relatively high ground motion, we selected a day during O4a with similar ground motion amplitude, and also plotted the corresponding glitches, represented by pink squares. There are fewer glitches in O4 between 18Hz and 30 Hz due to the scattering mitigation [soni2020reducing], but there is a higher rate in O4 at lower frequencies. While some high-amplitude glitches between 10 and 13 Hz were detectable during O3, most of the lower-frequency glitches observed during O4a would not have been “visible” during O3, given the sensitivity limitations of the interferometer at that time.
Figure 2: Characteristic strains during O3 in blue and O4a in pink, measured under similar microseismic motion amplitudes (or velocities). The blue circles highlight the impact of scattering glitches on the 18–30 Hz frequency band during O3, which was significantly reduced [soni2020reducing] in O4a (pink squares).
@@
As mentioned earlier, high ground motion (or elevated ground velocities) was the primary cause of scattering transients during O3. One class of these transients, referred to as Scattered Light or slow scattering [soni2020reducing, accadia2010noise], was characterized by arch-like features in their spectrograms. Given this, and based on the seasonal pattern in which ground motion in specific frequency bands tends to exhibit higher amplitudes during the winter months, we identified a potential correlation between such low-frequency glitches observed during O4a and elevated ground motion levels [alog68208]. This behavior is further supported by the ground motion data shown later in Figure 11.
@@
We analyze the Root Mean Square (RMS) of the ground motion within specific frequency ranges, known as Band-Limited RMS (BLRMS) [daw2004long]. The frequency bands considered correspond to the microseismic band-usually associated with ocean waves-and the anthropogenic band, which is generally linked to human activities such as passing trains or nearby vehicular traffic. The microseismic band is further subdivided into two distinct sub-bands. In this context, the analysis focuses on the following:
@@
Figure 5(b) presents the clusters identified via Agglomerative Clustering, using the optimal configuration determined by the Silhouette Score. Both techniques were implemented with the Scikit-learn library [pedregosa2011scikit]. As expected, the clustering process revealed three main groups, referred to here as Group L1 (purple), Group L2 (green), and Group L3 (orange).
An important advantage of this method is its ability to analyze the temporal evolution of glitches. Figure 6 presents the t-SNE output filtered month by month during O4a, providing an overview of which group was more prominent in each month. As mentioned, we selected the same number of glitches per month, enabling a balanced temporal analysis and preventing the dominance of months with higher glitch rates. Although this approach introduces an artificial uniformity across months, it allows for a clearer comparison of cluster behavior over time at this scale. If we had instead selected a total of 13,500 glitches randomly, without monthly filtering, months with higher glitch rates (see Figure 3) would have dominated the sample. A more detailed discussion of this selection strategy is provided in [ferreira2024using].
@@
The numbers displayed below each scatter plot represent the percentage contribution of each group for the respective month. Comparing May, June, and July reveals a high similarity in cluster distribution, with Group L2 (green) being the most prominent during all three months. The gray contours behind the data points indicate the estimated density, computed using Kernel Density Estimation (KDE) as implemented in Seaborn [waskom2021seaborn], where darker regions correspond to areas with higher data point concentrations.
@@
Group L2 primarily consists of glitch classes identified by Gravity Spy [wu2025advancing] as Extremely Loud, Blip, Koi Fish, Tomte, and others (examples are given in Figure 9). These glitches do not have a significant seasonal pattern (at least not as prominently as those in Group L1 and Group L3), which results in Group L2 gradually losing structure in the t-SNE output over time and being more dominant at the beginning of O4a, when the other groups were still less prominent. To further subdivide Group L2 into its specific subclasses, it would be necessary to rerun t-SNE on this subset. However, as previously mentioned, the primary objective of this study is to understand the behavior of the most frequent glitches that most affected the detector during O4a and to investigate potential correlations with environmental sensors surrounding the interferometer.
@@
As shown by the distribution of data points, the glitch rates for Group L1 (purple) and Group L3 (orange) exhibit a positive correlation with increasing amplitudes in the LMS and HMS bands (Figures 12(a) and 12(b)). In contrast, this trend is not observed for the anthropogenic band, suggesting no significant correlation in that case (Figures 12(c)). During the O4a period, some anthropogenic activities, such as logging [alog65597, alog65627] took place near the interferometer, leading to unusually high amplitudes in this band (exceeding $600\text{\,}\mathrm{n}\mathrm{m}\mathrm{/}\mathrm{s}$); this effect was observed, for example, in June, as illustrated by the variations shown in Figure 10(a). Coincidentally, microseismic noise was low at that time, and since most glitches were associated with microseismic motion, the overall glitch rate remained low. As a result, the region on the right side of Figure 12(c) behaves as an outlier when compared to the main distribution. This observation reinforces the conclusion that high amplitudes in the anthropogenic band do not significantly affect the glitch rate, confirming that this band did not contribute meaningfully to glitches with SNR >> 7.5, as also suggested in [alog77176].
@@
We quantified the correlation between each identified group and each ground-motion band using the Pearson correlation coefficient (rr[benesty2009pearson], which measures the strength of the linear relationship between two variables—in this case, the glitch rate and ground motion in different bands. Table 1 summarizes the correlation coefficients for each case, computed from more than 2,000 valid data points. Overall, the results confirm a strong correlation between Groups L1 and L3 and both the LMS and HMS bands, indicating a clear dependence of their occurrence on ground motion; their corresponding $p$-values (where smaller values mean the correlation is less likely to be due to chance) were below $10^{-100}$. Group L2 exhibits a weak correlation with the LMS band, which may partially arise from classification overlaps or a subclass with such dependence. As expected, none of the groups shows a significant correlation with the anthropogenic band. However, due to high-amplitude outliers associated with anthropogenic activities (previously mentioned), a small negative correlation appears for all three groups, with $p$-values equal to or below $0.05$. Despite these small $p$-values (likely driven by the large sample size), the corresponding correlation coefficients are very close to zero when considering only ground-motion amplitudes below $600$ nm/s (see Table 1), indicating no meaningful linear relationship.
@@
In conclusion, applying t-SNE to unclustered Omicron data from LLO enabled the identification of two glitch groups that follow the seasonal behavior of microseismic ground motion, providing strong support for the discussion reported in [soni2025ligo].

### v2

@@ 1 Introduction
Gravitational waves, predicted by Einstein’s general theory of relativity, have revolutionized our understanding of astrophysical phenomena by providing a new observational window into the Universe. Since the first direct detection of gravitational waves in 2015 by the Laser Interferometer Gravitational-Wave Observatory (LIGO) and the Virgo collaboration [1], ground-based detectors have been actively searching for signals during dedicated observing runs. Together with Virgo in Italy [2] and KAGRA in Japan [3], LIGO forms the LVK collaboration, which brings together a large community of researchers working toward the common goal of detecting an increasing number of gravitational-wave events.
@@
LIGO consists of two observatories designed to detect these spacetime ripples: one located in Hanford, Washington (LHO), and the other in Livingston, Louisiana (LLO). During gravitational wave searches, observatories operate in observing mode, and between observing runs, planned pauses are implemented to upgrade instrumentation and improve sensitivity [4]. The first detection occurred during the first observing run (O1), which lasted from September 12, 2015, to January 2016. On January 16, 2024, LIGO concluded the first part of its fourth observing run (O4a), which had started on May 24, 2023. As registered in the Gravitational-Wave Candidate Event Database (GraceDB) [5], there are already more than two hundred significant detection candidates identified during O4 produced by the mergers of compact objects such as black holes and neutron stars.
@@
LIGO also records various noise transients arising from instrumental or environmental factors, including ground motion, wind, and other disturbances [6, 7]. These non-astrophysical transient signals are commonly referred to as glitches. Glitches are non-stationary artifacts that contribute to the background in gravitational wave searches [8, 9, 10] and, in some cases, can closely mimic signals from compact binary coalescences (CBC). Efforts within CBC searches to mitigate this challenge are therefore essential and rely on different pipelines, such as PyCBC [11, 12], GstLAL [13, 14], MBTA [15, 16], and SPIIR [17, 18, 19], which incorporate likelihood weighting and other strategies to reduce the impact of glitches on search sensitivity.
@@
Glitch morphologies are typically analyzed using spectrograms [20], which facilitate their classification into distinct glitch groups. Different tools have been developed to analyze these unwanted signals. BayesWave [21, 22] and AWaRe [23], for instance, employ distinct approaches to reconstruct gravitational-wave signals contaminated by glitches. gengli [24, 25] is an open-source glitch generator useful to support a variety of testing and validation tasks, while other tools focus on glitch classification, improving classifier performance, or distinguishing true astrophysical signals from spurious noise transients [26, 27, 28, 29, 30]. One example is Gravity Spy, an interdisciplinary project that integrates machine learning, citizen science, and input from LIGO scientists to classify glitches [31, 32]. Glitches are named based on their morphology and/or known or suspected physical origin. Examples of their spectrogram morphologies can be found in [32, 8, 33, 34].
@@
The main motivation behind these efforts is to support the overarching goal of glitch studies: identifying their physical origin to eliminate or mitigate their impact on gravitational-wave detection. This work presents an analysis of the main glitch groups observed during O4a, investigating their potential correlations with environmental and/or instrumental variations using the methodology detailed in [35]. This investigation identifies patterns among glitches using the information provided by Omicron[36, 37], the primary tool in LIGO used to detect noise transients in detector data. When excess power is identified in the data, Omicron characterizes it by analyzing its time-frequency tiles in spectrograms produced from the $Q$-transform[38]. This information is stored in unclustered files, and after preparing these data, we employ t-distributed Stochastic Neighbor Embedding (t-SNE) [39], a dimensionality reduction technique, to project the high-dimensional output of Omicron (which includes numerous tile-based features) into a two-dimensional space.
@@
As an unsupervised technique, t-SNE does not require prior knowledge or a training dataset. Instead, it relies on probability-based similarity between data points to preserve local relative distances in the lower-dimensional representation. This makes t-SNE particularly useful for uncovering patterns in data without predefined assumptions. The resulting projections often reveal meaningful structures; as shown in [35, 40], cross-checking the clusters obtained from t-SNE with Gravity Spy labels demonstrates that, when applied to unclustered Omicron information, t-SNE produces groups that closely correspond to known glitch classes.
@@
In this work, the primary objective is to track the temporal evolution of glitches during the O4a observing run. To facilitate this analysis, we first apply t-SNE to project the high-dimensional data into a two-dimensional space and then use Agglomerative Clustering [41] to group similar glitches and assign labels. The Silhouette Score [42] is used to evaluate clustering quality and determine the optimal number of groups. These group labels serve primarily as a means to organize and visualize the temporal behavior of distinct glitch populations, to support their possible association with environmental or instrumental conditions, and to enable the classification of previously unseen data. This approach is entirely unsupervised and based solely on features derived from Omicron triggers. Further details about the method can be found in [35].
@@ 2 Glitches during O4a at LLO
The rate of glitches with a signal-to-noise ratio (SNR) exceeding 10 at LLO was higher during O4a (approximately 31 glitches per hour) compared to the third observing run, O3 (approximately 26 glitches per hour)[7]. The glitch rate during O3 was significantly influenced by the prevalence of scattering glitches, which dominated the transient population in that period. Scattering glitches occur when stray light couples back into the main beam after interacting with surfaces, whose identification is particularly challenging. Substantial efforts during O3 successfully identified and mitigated some of these scattering surfaces, resulting in a considerable reduction in the occurrence of such glitches[43, 44].
Despite these mitigation efforts, the transient rate during O4a surpassed that of O3; this increase was primarily driven by glitches occurring at low frequencies. Figure 1 illustrates the distribution of glitch frequencies below 100 Hz at LLO during O4a, highlighting this trend. The apparent gap in glitches around $14\text{\,}\mathrm{H}\mathrm{z}$ is due to masking by noise from the test masses’ roll modes and calibration lines at those frequencies. Additional peaks are also visible in the histogram, notably around 32, 39, 49, 61, 75, and 93 Hz. We compare the GravitySpy classes [32] in each of these frequency bins, and spectrogram analysis shows that the Tomte class accounts for more than 50% of glitches with peak frequencies near 32 and 39 Hz, while the 49 Hz peak consists of approximately 50% Blip Low Frequency. At 61 Hz, the dominant class is Blip Low Frequency; the peak around 75 Hz is primarily associated with Koi Fish, and the 93 Hz region includes a mix of Koi Fish and other classes. Some classes can share similar peak frequencies, but due to the complex nature of glitches, peak frequency alone is usually insufficient to distinguish between certain groups. This underscores the importance of analyzing glitches based on their morphology.
@@
The increase in glitch occurrence at lower frequencies is associated with the improved sensitivity of the interferometer in this frequency range [7]. Figure 2 illustrates this by comparing the characteristic strain at low frequencies (10–30 Hz) during O3, shown in blue, with that of O4a, shown in pink. This comparison highlights the enhanced sensitivity of the interferometer during O4a, which resulted in a higher detection rate of glitches (and of gravitational waves) within this frequency range. The blue dots represent the amplitude of scattering glitches over an eight-hour period on December 27, 2019 (during O3). Since these glitches were associated with relatively high ground motion, we selected a day during O4a with similar ground motion amplitude, and also plotted the corresponding glitches, represented by pink squares. There are fewer glitches in O4 between 18Hz and 30 Hz due to the scattering mitigation [43], but there is a higher rate in O4 at lower frequencies. While some high-amplitude glitches between 10 and 13 Hz were detectable during O3, most of the lower-frequency glitches observed during O4a would not have been “visible” during O3, given the sensitivity limitations of the interferometer at that time.
Figure 2: Characteristic strains during O3 in blue and O4a in pink, measured under similar microseismic motion amplitudes (or velocities). The blue circles highlight the impact of scattering glitches on the 18–30 Hz frequency band during O3, which was significantly reduced [43] in O4a (pink squares).
@@
As mentioned earlier, high ground motion (or elevated ground velocities) was the primary cause of scattering transients during O3. One class of these transients, referred to as Scattered Light or slow scattering [43, 45], was characterized by arch-like features in their spectrograms. Given this, and based on the seasonal pattern in which ground motion in specific frequency bands tends to exhibit higher amplitudes during the winter months, we identified a potential correlation between such low-frequency glitches observed during O4a and elevated ground motion levels [46]. This behavior is further supported by the ground motion data shown later in Figure 11.
@@
We analyze the Root Mean Square (RMS) of the ground motion within specific frequency ranges, known as Band-Limited RMS (BLRMS) [47]. The frequency bands considered correspond to the microseismic band-usually associated with ocean waves-and the anthropogenic band, which is generally linked to human activities such as passing trains or nearby vehicular traffic. The microseismic band is further subdivided into two distinct sub-bands. In this context, the analysis focuses on the following:
@@
Figure 5(b) presents the clusters identified via Agglomerative Clustering, using the optimal configuration determined by the Silhouette Score. Both techniques were implemented with the Scikit-learn library [48]. As expected, the clustering process revealed three main groups, referred to here as Group L1 (purple), Group L2 (green), and Group L3 (orange).
An important advantage of this method is its ability to analyze the temporal evolution of glitches. Figure 6 presents the t-SNE output filtered month by month during O4a, providing an overview of which group was more prominent in each month. As mentioned, we selected the same number of glitches per month, enabling a balanced temporal analysis and preventing the dominance of months with higher glitch rates. Although this approach introduces an artificial uniformity across months, it allows for a clearer comparison of cluster behavior over time at this scale. If we had instead selected a total of 13,500 glitches randomly, without monthly filtering, months with higher glitch rates (see Figure 3) would have dominated the sample. A more detailed discussion of this selection strategy is provided in [35].
@@
The numbers displayed below each scatter plot represent the percentage contribution of each group for the respective month. Comparing May, June, and July reveals a high similarity in cluster distribution, with Group L2 (green) being the most prominent during all three months. The gray contours behind the data points indicate the estimated density, computed using Kernel Density Estimation (KDE) as implemented in Seaborn [49], where darker regions correspond to areas with higher data point concentrations.
@@
Group L2 primarily consists of glitch classes identified by Gravity Spy [50] as Extremely Loud, Blip, Koi Fish, Tomte, and others (examples are given in Figure 9). These glitches do not have a significant seasonal pattern (at least not as prominently as those in Group L1 and Group L3), which results in Group L2 gradually losing structure in the t-SNE output over time and being more dominant at the beginning of O4a, when the other groups were still less prominent. To further subdivide Group L2 into its specific subclasses, it would be necessary to rerun t-SNE on this subset. However, as previously mentioned, the primary objective of this study is to understand the behavior of the most frequent glitches that most affected the detector during O4a and to investigate potential correlations with environmental sensors surrounding the interferometer.
@@
As shown by the distribution of data points, the glitch rates for Group L1 (purple) and Group L3 (orange) exhibit a positive correlation with increasing amplitudes in the LMS and HMS bands (Figures 12(a) and 12(b)). In contrast, this trend is not observed for the anthropogenic band, suggesting no significant correlation in that case (Figures 12(c)). During the O4a period, some anthropogenic activities, such as logging [51, 52] took place near the interferometer, leading to unusually high amplitudes in this band (exceeding $600\text{\,}\mathrm{n}\mathrm{m}\mathrm{/}\mathrm{s}$); this effect was observed, for example, in June, as illustrated by the variations shown in Figure 10(a). Coincidentally, microseismic noise was low at that time, and since most glitches were associated with microseismic motion, the overall glitch rate remained low. As a result, the region on the right side of Figure 12(c) behaves as an outlier when compared to the main distribution. This observation reinforces the conclusion that high amplitudes in the anthropogenic band do not significantly affect the glitch rate, confirming that this band did not contribute meaningfully to glitches with SNR >> 7.5, as also suggested in [53].
@@
We quantified the correlation between each identified group and each ground-motion band using the Pearson correlation coefficient (rr[54], which measures the strength of the linear relationship between two variables—in this case, the glitch rate and ground motion in different bands. Table 1 summarizes the correlation coefficients for each case, computed from more than 2,000 valid data points. Overall, the results confirm a strong correlation between Groups L1 and L3 and both the LMS and HMS bands, indicating a clear dependence of their occurrence on ground motion; their corresponding $p$-values (where smaller values mean the correlation is less likely to be due to chance) were below $10^{-100}$. Group L2 exhibits a weak correlation with the LMS band, which may partially arise from classification overlaps or a subclass with such dependence. As expected, none of the groups shows a significant correlation with the anthropogenic band. However, due to high-amplitude outliers associated with anthropogenic activities (previously mentioned), a small negative correlation appears for all three groups, with $p$-values equal to or below $0.05$. Despite these small $p$-values (likely driven by the large sample size), the corresponding correlation coefficients are very close to zero when considering only ground-motion amplitudes below $600$ nm/s (see Table 1), indicating no meaningful linear relationship.
@@
In conclusion, applying t-SNE to unclustered Omicron data from LLO enabled the identification of two glitch groups that follow the seasonal behavior of microseismic ground motion, providing strong support for the discussion reported in [7].